Diewold Company has two departments, Milling and Assembly. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Milling Department bases its rate on machine-hours, and the Assembly Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates:

   

Department
Milling Assembly
Direct labor-hours 7,600 78,000
Machine-hours 50,200 1,700
Total fixed manufacturing overhead cost $ 360,000 $ 548,000
Variable manufacturing overhead per machine-hour $ 3.00 -
Variable manufacturing overhead per direct labor-hour - $ 4.75

Required:

1. Compute the predetermined overhead rate to be used in each department. (Round your answers to 2 decimal places.)

        

2. Assume that the overhead rates you computed in (1) above are in effect. The job cost sheet for Job 407, which was started and completed during the year, showed the following:

    

Department
Milling Assembly
Direct labor-hours 4 12
Machine-hours 89 6
Materials requisitioned $ 750 $ 360
Direct labor cost $ 41 $ 180

   

Compute the total manufacturing cost assigned to Job 407. (Round your intermediate and final answers to 2 decimal places.)

         

3. Would you expect substantially different amounts of overhead cost to be charged to some jobs if the company used a plantwide overhead rate based on direct labor-hours instead of using departmental rates?

    

Yes
No
Answer #1

1) Molding overhead rate = (360000/50200)+3 = 10.17 per MH

Assembly overhead rate = (548000/78000)+4.75 = 11.78 per DLH

2) Job Cost

Direct material (750+360) 1110
Direct labor (41+180) 221
Molding overhead (89*10.17) 905.13
Assembly overhead (12*11.78) 141.36
Total 2377.49

3) Yes